MONT 108N – Mathematics Through Time Paper 1 – September 3, 2010

General Information

Your first formal writing assignment for the course will be an essay responding to one of the groups of questions below about our Natural World Cluster common reading, The Immortal Life of Henrietta Lacks, or if you prefer, to a question or questions of your choosing about the reading. Your paper should be prepared using MS Word or equivalent software, about 3 double-spaced pages in length (but don’t worry if you go a bit over or are a bit short – the length is not the main point). You will be saving the paper as a .doc file and submitting it by email to me by email to:

                        little@mathcs.holycross.edu

The due date/time is 5:00 p.m. on Friday, September 10.

Description

This is not a “research paper” – you need not consult sources other than The Immortal Life of Henrietta Lacks. Instead, you should think of your paper as a detailed defense of a particular conclusion about, or response to, some group of the questions in one of the topics below, written in a reasonably formal style incorporating correct grammar, spelling, and usage.

The specific questions you are addressing and your response(s) to them should be clearly stated in the first paragraph. Then the body of the paper should be a presentation of the reasoning and the evidence that lead you to your conclusion. Your job is to convince your reader of the soundness of your conclusion, starting “from the beginning.” That is, you should not assume the reader has been present for any of the discussions that we have had in class, or that you might have had with other Natural World Cluster students. You might try explaining possible opposing points of view, then giving your reasons for rejecting them, as part of your argument.

I will expect to see specific details from the text cited as evidence to support your assertions, and in many cases short quotations from the text will be the most effective way to provide evidence. Please identify any such quotations by the page(s) on which they appear. Use the following style: But Johns Hopkins Hospital was founded as an institution designed to help the less fortunate in society. Skloot states that “[i]t was built in 1889 as a charity hospital for the sick and poor, and it covered more than a dozen acres where a cemetery and insane asylum once sat in East Baltimore” (Skloot, p. 15). If you do use other sources and use quotations or other information, you should:

1) Identify direct quotations with quotes and a note in parentheses giving information where the quote can be found within your source,

2) List all your sources in a References section at the end of your paper:

a) For books, give the author, the title, the publisher and year of publication. b) For magazine or newspaper articles, give the author, the title of the magazine or newspaper, the date of publication, and the starting and final page of the specific article.

c) For web sites, give an author (if you can determine that), the full URL (web address) and the date you accessed it (since information on the web often changes!)

If you want to design a topic of your own, the same general ground rules apply.

A Suggestion

I will be happy to discuss (or read a first draft of) your paper and give you some preliminary comments by email. Or, you can come by my office hours if you want to “run your paper by me.” Alternatively, I think you may find it very helpful to have a first draft of your paper read by another student in the class. I can set up “reading pairs” if you are interested.

Evaluation

I will provide written comments on your work, and assign two grades for each paper – one for how well your conclusions are presented and supported (in other words, for how convincing your arguments are), the other for how well your writing follows the standards for formal written English. (For instance, how well is the paper subdivided into paragraphs each addressing a particular item in your overall argument? Are the paragraphs arranged in a logical, recognizable sequence? Are the sentences within each paragraph ordered well? Are they grammatically correct? Are there awkward sentences? Are there overly flowery, overly colloquial, or incorrectly-used words or phrases? Is punctuation used correctly? Are there spelling and/or typographical errors?)

Everyone will have the opportunity to submit a revised version of the paper after receiving my comments. If you choose to resubmit, I’ll evaluate the new version in the same way and count only the better of the two sets of marks. The due date for resubmissions will be announced when the papers are returned.

Topic 1 – What’s in a Name?

One of the possibly puzzling aspects of the story told in The Immortal Life of Henrietta Lacks is the way Richard TeLinde, George Gey and others seem to have systematically deceived the world about the name of the woman from whom the HeLa cells were collected. From the evidence in the book, what do you think the reasons were for their attitudes and actions? Were they trying to hide something? Were they ashamed of something they had done? Were they trying to maintain control over use of the cells? Were they concerned about having to share the money that the cells were generating? What would have changed if the world knew that HeLa came from a woman named Henrietta Lacks?

One of the fundamental ethical principles of current medical research involving human subjects is that those subjects should understand and give their consent to all treatments and procedures that are performed on them. The next two groups of questions deal with different aspects of how the idea of informed consent enters into the story of The Immortal Life of Henrietta Lacks.

Topic 2 – What Were the Ethical Standards of the Time?

At the time Henrietta Lacks was treated at Johns Hopkins, there were no laws on the books in the U.S. requiring doctors to get consent from patients before performing most treatments or tests. From the evidence in The Immortal Life of Henrietta Lacks, do you think that the doctors who treated her believed that they were acting according to high ethical standards? Do you think Henrietta was treated reasonably? Be sure to distinguish whether you are answering this according to the standards of the time or according to your own standards. Were those standards of the time reasonable? Were there studies or tests described in the book that were done in the U.S. at roughly the same time that were clearly abuses of the rights of the subjects? Are there times when what is legal and what is ethical conflict?

Topic 3 – Should Doctors Have to Educate Patients to Get Consent?

Think about the situation you would be in as a doctor if you had a patient like Henrietta (or Elsie) Lacks and you were able to see the potential for gaining knowledge or a developing a new treatment that would help others by doing some test or procedure. What does it mean to get informed consent if the person involved does not (or can not) understand the basic science involved? Do you think the researcher should have the obligation to educate the subject so that they can understand? Is that even always possible? Are there times when the possibility of helping many others outweighs the rights of subjects? And is this an issue for research with tissue such as embryonic stem cells? (For this topic, you might want to consult other sources, to see, for instance what the Nuremberg Code, the Helsinki Declaration, and the U.S. Belmont Report actually say about these issues.)